A ball mass of 0.2kg strikes on inclined surface and makes angle 60 degree and speed change from 30m/s to 20m/s what is the magnitude of impulse?


1
Expert's answer
2022-04-26T14:55:19-0400

"∆p=mv_1+mv_2\cos60°=8~\frac{kg\cdot m}s ."